OpenStudy (anonymous):

Write equation of the line in slope intercept form that passes through the two points (-2,1) (2,4 )

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

slope=3/4 so equation=y-1=3/4(x+2) solving you get 4y=3x+10

OpenStudy (anonymous):

@Jessica_Love :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@Jessica_Love u get it,right?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

wait would i divide by 4

OpenStudy (anonymous):

@Jessica_Love ya since you want it in slope intercept form..:)

OpenStudy (anonymous):

kk